A.     Solve the following problems. Write your answer on a sheet of paper.


1.      Given that sin α =  and sin β = , find 𝑡𝑎𝑛 (α + β) if both α and β are in QIV.

2.      Given that csc A = , A in QI, and sec B = , sin B < 0, find

a.      cos (A – B)

b.     tan (A – B)

3.      If tan (x + y) =  and tan y = , what is tan x?

4.      The point (9, –5) lies on the terminal side of the angle 𝜃 in standard position. Find (sin 𝜃 + cos 𝜃).


B.      Solve the following problems. Write your answer on a sheet of paper.


1.      If cos t = , what is cos 2t?

2.      Use half – angle identities to find the exact value of tan 22.50 and sin 150.


1
Expert's answer
2021-03-08T03:08:16-0500

1.


"\sin \alpha=a, -1\leq a\leq0"

"\sin\beta=b, -1\leq a\leq0"

"\cos \alpha=\sqrt{1-\sin^2 \alpha}=\sqrt{1-a^2}"

"\cos\beta =\sqrt{1-\sin^2 \beta}=\sqrt{1-b^2}"


"\tan(\alpha+\beta)=\dfrac{\tan\alpha+\tan\beta}{1-\tan\alpha\tan\beta}"

"=\dfrac{\dfrac{a}{\sqrt{1-a^2}}+\dfrac{b}{\sqrt{1-b^2}}}{1-\dfrac{a}{\sqrt{1-a^2}}\dfrac{b}{\sqrt{1-b^2}}}"

"=\dfrac{a\sqrt{1-b^2}+b\sqrt{1-a^2}}{\sqrt{1-a^2}\sqrt{1-b^2}-ab}"

"\tan(\alpha+\beta)=\dfrac{a\sqrt{1-b^2}+b\sqrt{1-a^2}}{\sqrt{1-a^2}\sqrt{1-b^2}-ab}"

2.


"\csc A=\dfrac{1}{\sin A}=a, a\geq 1"


"\sin A=\dfrac{1}{a},"

"\cos A=\sqrt{1-\sin^2 A}=\sqrt{1-(\dfrac{1}{a})^2}=\dfrac{\sqrt{a^2-1}}{a}"

"\sec B=\dfrac{1}{\cos B}=b, |b|\geq 1"

"\cos B=\dfrac{1}{b}"

"\sin B=-\sqrt{1-\cos^2 B}=-\sqrt{1-(\dfrac{1}{b})^2}=-\dfrac{\sqrt{b^2-1}}{|b|}"

a.


"\cos(A-B)=\cos A\cos B+\sin A\sin B"

"=\dfrac{\sqrt{a^2-1}}{a}(\dfrac{1}{b})+\dfrac{1}{a}(-\dfrac{\sqrt{b^2-1}}{|b|})"

"b\geq1, a\geq1:\cos(A-B)=\dfrac{\sqrt{a^2-1}-\sqrt{b^2-1}}{ab}"


"b\leq-1, a\geq1:\cos(A-B)=\dfrac{\sqrt{a^2-1}+\sqrt{b^2-1}}{ab}"

b.


"\tan(A-B)=\dfrac{\sin(A-B)}{\cos(a-B)}"

"\sin(A-B)=\sin A\cos B-\cos A\sin B"

"=\dfrac{1}{a}(\dfrac{1}{b})-\dfrac{\sqrt{a^2-1}}{a}(-\dfrac{\sqrt{b^2-1}}{|b|})"

"b\geq1, a\geq1:\cos(A-B)=\dfrac{\sqrt{a^2-1}-\sqrt{b^2-1}}{ab}"


"\sin(A-B)=\dfrac{1+\sqrt{(a^2-1)(b^2-1)}}{ab}"

"\tan(A-B)=\dfrac{1+\sqrt{(a^2-1)(b^2-1)}}{\sqrt{a^2-1}-\sqrt{b^2-1}}"


"b\leq-1, a\geq1:\cos(A-B)=\dfrac{\sqrt{a^2-1}+\sqrt{b^2-1}}{ab}"


"\sin(A-B)=\dfrac{1-\sqrt{(a^2-1)(b^2-1)}}{ab}"

"\tan(A-B)=\dfrac{1-\sqrt{(a^2-1)(b^2-1)}}{\sqrt{a^2-1}+\sqrt{b^2-1}}"

3.


"\tan(x+y)=\dfrac{\tan x+\tan y}{1-\tan x \tan y}"

"\tan (x+y)-\tan x \tan (x+y)\tan y=\tan x+\tan y"

"\tan x=\dfrac{\tan (x+y)-\tan y}{1+\tan (x+y) \tan y}"

Given "\tan(x+y)=a, \tan y=b"


"\tan x=\dfrac{a-b}{1+ab}"

4.


"\sqrt{(9)^2+(-5)^2}=\sqrt{106}"

"\sin \theta=\dfrac{-5}{\sqrt{106}}, \cos\theta=\dfrac{9}{\sqrt{106}}"

"\sin \theta+\cos \theta =-\dfrac{5}{\sqrt{106}}+\dfrac{9}{\sqrt{106}}=\dfrac{4}{\sqrt{106}}"

"\sin \theta+\cos \theta =\dfrac{4}{\sqrt{106}}"

B.

1.     


"\cos2t=2\cos^2t-1"

Given "\cos t=a"


"\cos2t=2a^2-1"

2.


"\sin 15\degree=\sqrt{\dfrac{1-\cos 30\degree}{2}}=\sqrt{\dfrac{1-\dfrac{\sqrt{3}}{2}}{2}}"

"\sin 15\degree=\dfrac{\sqrt{2-\sqrt{3}}}{2}"

"\tan 22.5\degree=\sqrt{\dfrac{1-\cos 45\degree}{1+\cos45\degree}}=\sqrt{\dfrac{1-\dfrac{\sqrt{2}}{2}}{1+\dfrac{\sqrt{2}}{2}}}"

"=\sqrt{\dfrac{2-\sqrt{2}}{2+\sqrt{2}}}=\dfrac{2-\sqrt{2}}{\sqrt{2}}=\sqrt{2}-1"


"\tan 22.5\degree=\sqrt{2}-1"